# An Open-Source Annotation System for Education Transcript Data

### v1

@@ Abstract
Guanzhong Pan1, Mei Tan2, Hyunji Nam2, Lucía Langlois2, James Malamut2, Liliana Deonizio2, Dorottya Demszky2 1Carnegie Mellon University, 2Stanford University
@@ 1 Introduction
First, no purpose-built, adaptive tool exists for annotating educational transcripts with iterative co-design and calibration support. Most existing platforms target general NLP taggings like named entity recognition (NER) or part-of-speech (POS) tagging, and lack support for annotating high-inference pedagogical features such as the discourse moves of teachers and students–and the latent cognitive processes they enact (Nakayama et al., 2018; Perry, 2021). Qualitative analysis platforms ATLAS.ti Scientific Software Development GmbH (2023); Dhakal (2022) support open coding but lack support for structured, utterance-level annotation and real-time analysis for collaborative calibration. As a result, many researchers rely on spreadsheets or ad hoc solutions, which are time-consuming to build, often clunky to use, and are poorly integrated with modern machine learning workflows.

### v2

@@ Abstract
Guanzhong Pan1, Mei Tan2, Hyunji Nam2, Lucía Langlois2, James Malamut2, Liliana Deonizio2, Dorottya Demszky2 1Carnegie Mellon University, 2Stanford University arthurp@andrew.cmu.edu, {mxtan, hjnam, lucilang, jmalamut, lilianas, ddemszky}@stanford.edu
@@ 1 Introduction
First, no purpose-built, adaptive tool exists for annotating educational transcripts with iterative co-design and calibration support. Most existing platforms target general NLP taggings like named entity recognition (NER) or part-of-speech (POS) tagging, and lack support for annotating high-inference pedagogical features such as the discourse moves of teachers and students–and the latent cognitive processes they enact (Nakayama et al., 2018; Honnibal and Montani, 2017; Perry, 2021). Qualitative analysis platforms ATLAS.ti Scientific Software Development GmbH (2023); Dhakal (2022) support open coding but lack support for structured, utterance-level annotation and real-time analysis for collaborative calibration. As a result, many researchers rely on spreadsheets or ad hoc solutions, which are time-consuming to build, often clunky to use, and are poorly integrated with modern machine learning workflows.